OpenStudy (anonymous):

What is the amplitude and period of f(t)=cos 3t

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

But for \(f(t) = \cos\left(3t\right)\) this is a cosine function so you there is an amplitude and a period. The amplitude is the absolute value of the coefficient on the outside of function, which is 1, so the amplitude is 1. The period is \(\frac{2\pi}{B}\) where B is the coefficient inside the parentheses, in this case, 3. So the period is \(\frac{2\pi}{3}\).